OpenStudy (brooke1556):

What is the slope-intercept form equation of the line that passes through (2, 4) and (4, 10)? y = −3x − 2 y = −3x + 2 y = 3x + 2 y = 3x − 2

OpenStudy (sloppycanada):

Do you know the equation of a line is y=mx+b Since we are given two sets of coordinates, we must find the slope. \[\frac{ y_2-y_1 }{ x_2-x_1 }\]

OpenStudy (sloppycanada):

So using the equation that was given, y2 - y1 =? x2 - x1 = ?

OpenStudy (brooke1556):



OpenStudy (sloppycanada):

Correct, can we simplify 6/2?

OpenStudy (brooke1556):

Yes, divide by two and it becomes 3?

OpenStudy (sloppycanada):

Bingo! We now know the slope now according to the equation y = mx+b m= slope So we're looking at the equation - y = 3x+b From here, all we do is substitute the coordinate points of (4,10) for x and y respectively.

OpenStudy (brooke1556):

What do I do with (4,10) ?